OpenStudy (xenesis98):

PLEASE !!!! for a medal Solve 2cos^2x + cosx − 1 = 0 for x over the interval [0, 2 π ). a.π and π/3 b.π, π/3, and 5π/3 c.1 and 2π/3 d.1, 2π/3, and 4π/3 e.1, π/3, and 5π/3

OpenStudy (anonymous):

\[2\cos ^2x+2\cos x-\cos x-1=0\] make factors and find value of cos x and then x

OpenStudy (xenesis98):

idk how

OpenStudy (xenesis98):

and i dont have time

OpenStudy (anonymous):

take cos x common from first two and -1 from last two

OpenStudy (xenesis98):

cos^2x + cosx = 0?

OpenStudy (anonymous):

\[2\cos x \left( \cos x+1 \right)-1\left( \cos x+1 \right)=0\] \[\left( \cos x+1 \right)\left( 2\cos x-1 \right)=0\] equate each factor to 0 and find value of x. will you show me ?

OpenStudy (anonymous):

either cos x+1=0, cos x=-1 can you tell at what angle value of cosx =-1

OpenStudy (xenesis98):

idk

OpenStudy (xenesis98):

x=π±2πn,π/3±2πn,5π/3±2πn

OpenStudy (xenesis98):

i got that thing for first thing you asked

OpenStudy (anonymous):

\[\cos x=-1=\cos \pi \in 0 \le x \le 2 \pi,x=\pi\]

OpenStudy (anonymous):

\[2 \cos x=1,\cos x=\frac{ 1 }{ 2 }=\cos \frac{ \pi }{ 3 },\cos \left( 2\pi-\frac{ \pi }{ 3 } \right)\] x=?

OpenStudy (xenesis98):

cos(5x/3)

OpenStudy (xenesis98):

wait i made a mistake

OpenStudy (xenesis98):

1/2